OpenStudy (anonymous):

Solve the system of equations. y = 7x + 43 y = 2x + 13 a. ( -6, 1) c. ( 6, 1) b. ( 6, -1) d. No solution

OpenStudy (anonymous):

@Gummibear127

OpenStudy (kidrah69):

plug in equation 1 into equation 2

OpenStudy (kidrah69):

\[(7x + 43)= 2x + 13\]

OpenStudy (kidrah69):

we want to isolate X so subtract 43 from both sides

OpenStudy (kidrah69):

13-43=?

OpenStudy (anonymous):

Hey it looks like @kidrah69 is already helping you and im not really good with this any way sorry

OpenStudy (anonymous):

it says -30

OpenStudy (anonymous):

i wanted her to help though

OpenStudy (kidrah69):

yep thats correct so now we have \[7x=2x-30\] add 2x to 7x 2+7=?

OpenStudy (kidrah69):

@eminem-101 if you want her to help i can stop :)

OpenStudy (anonymous):

9

OpenStudy (anonymous):

it is fine

OpenStudy (kidrah69):

ok good so we have 9x=-30 when you divide what do you get?

OpenStudy (anonymous):

0

OpenStudy (kidrah69):

No -30/9 is NOT 0

OpenStudy (anonymous):

oh it's -3.333

OpenStudy (radar):

Why did you add the 2x to the 7x ?? Be careful.

OpenStudy (kidrah69):

dont you combine like terms?

OpenStudy (kidrah69):

oh wait you would subtract

OpenStudy (kidrah69):

right so 7-2

OpenStudy (radar):

Yes if there is a + between them. You have an equal sign.

OpenStudy (anonymous):

5

OpenStudy (kidrah69):

right so -30/5=?

OpenStudy (anonymous):

-6

OpenStudy (kidrah69):

yep :) now plug that into the 2nd equation \[y=2(-6)+13\] solve that

OpenStudy (anonymous):

1

OpenStudy (kidrah69):

\[(7x+43)=2x+13 \] subtract 43 from both sides to isolate X \[7x=2x−30\] subtract 2 from 7 \[5x=-30\] divide x=-6 plug in -6 to equation 2 \[y=2(−6)+13\] y=-1

OpenStudy (kidrah69):

so whats the final answer and i rewrote that so you can see what we did :)

OpenStudy (anonymous):

i got 1 not -1

OpenStudy (kidrah69):

yeah i accidently put that negative =/ so whats the final answer ? We have (-6,1)

OpenStudy (anonymous):

(-6,1) is the answer right

OpenStudy (kidrah69):

yep :) xD

OpenStudy (anonymous):

that you

OpenStudy (kidrah69):

that i think you mean thanks you're welcome ;)

OpenStudy (anonymous):

sorry you are right

OpenStudy (kidrah69):

lol o need to be sorry :P